# On $f$-Divergence Principled Domain Adaptation: An Improved Framework

### v1

@@ Abstract
Unsupervised domain adaptation (UDA) plays a crucial role in addressing distribution shifts in machine learning. In this work, we improve the theoretical foundations of UDA proposed by Acuna et al. (2021) by refining their $f$-divergence-based discrepancy and additionally introducing a new measure, $f$-domain discrepancy ($f$-DD). By removing the absolute value function and incorporating a scaling parameter, $f$-DD yields novel target error and sample complexity bounds, allowing us to recover previous KL-based results and bridging the gap between algorithms and theory presented in Acuna et al. (2021). Leveraging a localization technique, we also develop a fast-rate generalization bound. Empirical results demonstrate the superior performance of $f$-DD-based domain learning algorithms over previous works in popular UDA benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Machine learning often faces the daunting challenge of domain shift, where the distribution of data in the testing environment differs from that used in training. Unsupervised Domain Adaptation (UDA) arises as a solution to this problem. In UDA, models are allowed to access to labelled source domain data and unlabelled target domain data, while the ultimate goal is to find a model that performs well on the target domain. The mainstream theoretical foundations of UDA, and more broadly, domain adaptation (Quionero-Candela et al., 2009), primarily rely on the seminal works of discrepancy-based theory (Ben-David et al., 2006, 2010). In particular, Ben-David et al. (2006, 2010) characterize the error gap between two domains using a hypothesis class-specified discrepancy measure e.g., $\mathcal{H}\Delta\mathcal{H}$-divergence. While these works initially focus on binary classification tasks and zero-one loss, Mansour et al. (2009a) extend the theory to a more general setting. Subsequently, this kind of theoretical framework was extended by various works (Cortes et al., 2015; Germain et al., 2013, 2016; Zhang et al., 2019, 2020; Acuna et al., 2021; Shui et al., 2022), all sharing some common properties such as the ability to estimate the proposed discrepancy from finite unlabeled samples. Importantly, these theoretical results often inspire the design of new algorithms, such as domain-adversarial training of neural networks (DANN) (Ganin et al., 2016) and Margin Disparity Discrepancy (MDD) (Zhang et al., 2019), directly motivated by $\mathcal{H}\Delta\mathcal{H}$-divergence.
@@
Recently, Acuna et al. (2021) propose an $f$-divergence-based domain learning framework, which generalizes various previous frameworks (e.g., those based on $\mathcal{H}\Delta\mathcal{H}$-divergence and total variation) and have demonstrated great empirical successes. However, we argue that this framework has potential limitations, at least in three aspects.
@@
First, their discrepancy measure is based on the variational representation of $f$-divergence in Nguyen et al. (2010) (cf. Lemma 2.1). Although this variational formula is commonly adopted, its weakness has been pointed out in several works (Ruderman et al., 2012; Jiao et al., 2017). For instance, from this formula, one cannot recover the well-known Donsker and Varadhan’s (DV) representation of KL divergence (Donsker & Varadhan, 1983). This reveals a second limitation: some existing domain adaptation or transfer learning theories are based on the DV representation of KL, such as (Wu et al., 2020; Bu et al., 2022; Wang & Mao, 2023a), and the framework by Acuna et al. (2021), although including KL as a special case of $f$-divergence, can not unify those theories. Furthermore and more critically, the discrepancy measure proposed by Acuna et al. (2021) contains an absolute value function, while the original variational representation does not. Notably, this absolute value function is necessary in their derivation of an target error upper bound, and the fundamental reason behind this is still due to the weak version of the variational representation relied upon. Specifically, a variational representation of an $f$-divergence is a lower bound of the divergence, but using a weak lower bound creates technical difficulties in proving an upper bound for target domain error. Acuna et al. (2021) chooses to add an absolute value function accordingly, potentially leading to an overestimation of the corresponding $f$-divergence. In fact, this absolute value function is removed in their proposed algorithm, termed $f$-Domain Adversarial Learning ($f$-DAL). While $f$-DAL outperforms both DANN and MDD in standard benchmarks, this choice exhibits a clear gap between their theory and algorithm.
@@
In this work, to overcome these limitations and explore the full potential of $f$-divergence, we present an improved framework for $f$-divergence-based domain learning theory. Specifically, we employ a more advanced variational representation of $f$-divergence (cf. Lemma 2.2), independently developed by Agrawal & Horel (2020, 2021) and Birrell et al. (2022). After introducing some preliminaries, the organization of the remainder of the paper and our main contributions are summarized below.
  • [leftmargin=.2cm,topsep=-0cm, parsep=-0cm,labelsep=-0.3cm,align=left]
@@
  • In Section 3, we revisit the theoretical analysis in Acuna et al. (2021), where we refine their $f$-divergence-based discrepancy by Lemma 2.2 while retaining the absolute value function in the definition. The resulting target error bound (cf. Lemma 3.1) and the KL-based generalization bound (cf. Theorem 3.1) complement the theoretical framework of Acuna et al. (2021).
@@
  • In Section 4, we design a novel $f$-divergence-based domain discrepancy measure, dubbed $f$-DD. Specifically, we eliminate the absolute value function from the definition used in Section 3, incorporating a scaling parameter instead. We then derive an upper bound for the target error (cf. Theorem 4.1) and the sample complexity bound for our $f$-DD. The generalization bound based on empirical $f$-DD naturally follows from these results. Notably, the obtained target error bound allows us to recover the previous KL-based result in Wang & Mao (2023a) (cf. Corollary 4.1).
@@
  • In Section 5, to improve the convergence rate of our $f$-DD based bound, we sharpen the bound using a localization technique (Bartlett et al., 2005; Zhang et al., 2020). The localized $f$-DD allows for a crisp application of the local Rademacher complexity based concentration results (Bartlett et al., 2005), while also enabling us to achieve a fast-rate target error bound (cf. Theorem 5.1). As a concrete example, we present a generalization bound based on localized KL-DD (cf. Theorem 5.2), where our proof techniques are directly connected to fast-rate PAC-Bayesian bounds (Catoni, 2007; Alquier, 2021) and fast-rate information-theoretic generalization bounds (Hellström & Durisi, 2021, 2022; Wang & Mao, 2023c).
@@
  • In Section 6, we conduct empirical studies based on our $f$-DD framework. We show $f$-DD outperforms the original $f$-DAL in three popular UDA benchmarks, with the best performance achieved by Jeffereys-DD. Additionally, we note that the training objective in $f$-DAL aligns more closely with our theory than with Acuna et al. (2021) (cf. Proposition 1). We also show that adding the absolute value function leads to overestimation and that optimizing the scaling parameter in $f$-DD may not be necessary in practical algorithms.
@@ 2 Preliminaries
Throughout the paper, unless otherwise noted, we will use a capitalized letter to denote a random variable, and use the corresponding lower-case letter to denote its realization. Let $\mathcal{X}$ and $\mathcal{Y}$ be the input space and the label space. Let $\mathcal{H}$ be the hypothesis space, where each $h\in\mathcal{H}$ is a hypothesis mapping $\mathcal{X}$ to ${\cal Y}$.
Consider a single-source domain adaptation setting, where $\mu$ and $\nu$ are two unknown distributions on $\mathcal{X}\times\mathcal{Y}$, characterizing respectively the source domain and the target domain. Let $\mathcal{S}=\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}{\sim}\mu^{\otimes n}$ be a labeled source-domain sample and $\mathcal{T}=\{X_{j}\}_{j=1}^{m}{\sim}\nu^{\otimes m}$ be an unlabelled target-domain sample. We use $\hat{\mu}$ and $\hat{\nu}$ to denote the empirical distributions on ${\cal X}$ corresponding to $\mathcal{S}$ and $\mathcal{T}$, respectively. The objective of UDA is to find a hypothesis $h\in\mathcal{H}$ based on $\mathcal{S}$ and $\mathcal{T}$ that “works well” on the target domain.
Let $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ be a symmetric loss (i.e. $\ell(y,y)=0$ for all $y\in{\cal Y}$). The population risk for each $h\in\mathcal{H}$ in the target domain (i.e. target error) is defined as $R_{\nu}(h)\triangleq{{\mathbb{E}_{(X,Y)\sim\nu}}\left[\ell(h(X),Y)\right]},$ and the population risk in the source domain, $R_{\mu}(h)$, is defined in the same way. Since $\mu$ and $\nu$ are unknown to the learner, one often uses recourse to the empirical risk in the source domain, which, for a given $\mathcal{S}$, is defined as $R_{\hat{\mu}}(h)\triangleq{{\mathbb{E}_{(X,Y)\sim\hat{\mu}}}\left[\ell(h(X),Y)\right]}={\frac{1}{n}\sum_{i=1}^{n}\ell(h(X_{i}),Y_{i})}.$
With a little abuse of the notation, we will simply use $\ell(h,h^{\prime})$ to represent $\ell(h(x),h^{\prime}(x))$ when the same $x$ is evaluated. Additionally, following the conventional literature on domain adaptation theory (Mansour et al., 2009a), we assume that the loss function satisfies the triangle property [^1] The triangle property of loss function indicates that $\ell(y_{1},y_{2})\leq\ell(y_{1},y_{3})+\ell(y_{3},y_{2})$ for any $y_{1},y_{2},y_{3}\in\mathcal{Y}$.. For the readers’ convenience, a summary of all the notations is provided in Table 5 in the Appendix.
@@
Let $P$ and $Q$ be two distributions on $\Theta$. Let ϕ:+:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}$ be a convex function with $\phi(1)=0$. If $P\ll Q$, then $f$-divergence is defined as
${\rm D}_{\phi}(P||Q)\triangleq{{\mathbb{E}_{Q}}\left[\phi\left(\frac{dP}{dQ}\right)\right]},$
where $\frac{dP}{dQ}$ is a Radon-Nikodym derivative.
The $f$-divergence family contains many popular divergences. For instance, letting $\phi(x)=x\log{x}$ (or $x\log{x}+c(x-1)$ for any constant $c$) recovers the definition of KL divergence.
The $f$-divergence discrepancy measure by Acuna et al. (2021) is motivated by the variational formula of $f$-divergence that utilizes the Legendre transformation (LT).
Lemma 2.1 (Nguyen et al. (2010)).
${\rm D}_{\phi}(P||Q)=\sup_{g\in\mathcal{G}}{{\mathbb{E}_{\theta\sim P}}\left[g(\theta)\right]}-{{\mathbb{E}_{\theta\sim Q}}\left[\phi^{*}(g(\theta))\right]},$
However, it is well-known that the variational formula in Lemma 2.1 could not recover the the Donsker and Varadhan’s (DV) representation of KL divergence (cf. Lemma B.1). We will elaborate on this later.
Recently, Birrell et al. (2022) and Agrawal & Horel (2020, 2021) concurrently introduce a novel variational representation for $f$-divergence, which is also implicitly stated in Ben-Tal & Teboulle (2007, Theorem 4.2), as given below.
Lemma 2.2 (Agrawal & Horel (2020, Corollary 3.5)).
Any $f$-divergence can be written as
This variational representation is a “shift transformation” of Lemma 2.1 (i.e., $g\to g+\alpha$). Note that this representation shares the same optimal solution as Lemma 2.1 (clearly identified as the corresponding $f$-divergence), but Lemma 2.2 is considered tighter in the sense that the representation in Lemma 2.2 is flatter around the optimal solution. Birrell et al. (2022) provides a comprehensive study to justify this, and they also show that Lemma 2.2 can accelerate numerical estimation of $f$-divergences.
Here, to illustrate the advantage of Lemma 2.2, we use KL divergence as an example. Specifically, let $\phi(x)=x\log{x}-x+1$, then its conjugate function is $\phi^{*}(y)=e^{y}-1$. Substituting $\phi^{*}$ into Lemma 2.1, we have the LT-based KL:
On the other hand, Lemma 2.2 will give us
$\mathrm{D_{KL}}\left(P||Q\right)=$ $\sup_{g\in\mathcal{G}}{{\mathbb{E}_{P}}\left[g(\theta)\right]}\!-\!\inf_{\alpha\in\mathbb{R}}\left\{{{\mathbb{E}_{Q}}\left[e^{g(\theta)+\alpha}\right]}\!-\!1\!-\!\alpha\right\}$
$=$ $\sup_{g\in\mathcal{G}}{{\mathbb{E}_{P}}\left[g(\theta)\right]}-\log{{\mathbb{E}_{Q}}\left[e^{g(\theta)}\right]},$ (2)
where the optimal $\alpha^{*}=-\log{{{{\mathbb{E}_{Q}}\left[e^{g(\theta)}\right]}}}$.
Notice that Eq. (2) immediately recovers the DV representation of KL. Since $\log(x)\leq x-1$ for $x>0$, it is evident that, as a lower bound of KL divergence, Eq. (2) is pointwise tighter than Eq. (1). In Appendix C, we also show the variational representations of some other divergences.
In the context of UDA, it may be tempting to think that using a point-wise smaller quantity (in Lemma 2.1), as the key component of an upper bound for target error, is essentially desired. However, neither Lemma 2.1 nor Lemma 2.2 is able to directly give such an upper bound. To elaborate, as $\phi^{*}(x)\geq x$ when $\phi(1)=0$ (cf. Lemma B.2), both Lemma 2.1 and Lemma 2.2 imply that $\mathrm{D}_{\phi}(P||Q)\leq\sup_{g}{{\mathbb{E}_{P}}\left[g(\theta)\right]}-{{\mathbb{E}_{Q}}\left[g(\theta)\right]}$. Bear this in mind, UDA typically requires an upper bound for the quantity $\sup_{h}{{\mathbb{E}_{\nu}}\left[\ell\circ h(X)\right]}-{{\mathbb{E}_{\mu}}\left[\ell\circ h(X)\right]}$, and simply restricting $\mathcal{G}$ in Lemma 2.1 and Lemma 2.2 to a composition of $\mathcal{H}$ and $\ell$ can only provide a lower bound for such a quantity.
Before we dedicate to an improved discrepancy measure, we first revisit the absolute discrepancy in Acuna et al. (2021) by using Lemma 2.2 instead. This also serves as a review of the common developments in the DA theory.
3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
Based on Lemma 2.2, we refine the discrepancy measure of Acuna et al. (2021, Definition 3) as follows.
Definition 3.1.
For a given $h\in\mathcal{H}$, the $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}$ discrepancy is
@@ 3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
Removing the absolute value function in $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)$ does not alter its non-negativity. To see this, consider $h^{\prime}=h$. By the definitions of $\phi$ and $\phi^{*}$, $\inf_{\alpha}\phi^{*}(\alpha)-\alpha=\phi(1)=0$, we have ${{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime})\right]}=I_{\nu}^{\phi}(h,h^{\prime})=0$. Consequently, since $h$ exists in $\mathcal{H}$, $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)\geq 0$ holds even without the absolute value function. In addition, due to the absolute value function, the relation between $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)$ with the one in Acuna et al. (2021) is no longer clear.
@@
$*.\displaystyle R_{\nu}(h)\leq R_{\mu}(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+\lambda^{*}.$
This error bound shares similarities with the previous works (Ben-David et al., 2006, 2010; Zhang et al., 2019, 2020; Acuna et al., 2021). For example, the third term $\lambda^{*}$ is the ideal joint risk for the DA problem. As widely discussed in prior studies, this term captures the inherent challenge in the DA task and might be inevitable (Ben-David et al., 2010; Ben-David & Urner, 2012).
In the sequel, we will give a Rademacher complexity based generalization bound for the target error. Let ^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ denote the empirical Rademacher complexity of function class =$\mathcal{F}=\{f:\mathcal{X}\to\mathbb{R}\}$ for some sample $S$ (Bartlett & Mendelson, 2002). Notice that a shift transformation of a function class will not change its Rademacher complexity, so the generalization bound based on $\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}$ closely resembles the one presented in Acuna et al. (2021, Theorem 3), which contains a Lipschitz constant of $\phi^{*}$. Here, we give a generalization bound specialized for $\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}$, wherein the Lipschitz constant of $\phi^{*}$ can be explicitly determined.
@@
$R_{\nu}(h)\leq R_{\hat{\mu}}$ $(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\hat{\mu}||\hat{\nu})+2e\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})+4\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})$
$+{\lambda}^{*}+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}+\sqrt{\frac{\log(1/\delta)}{m}}\right),$

### v2

@@ Abstract
Unsupervised domain adaptation (UDA) plays a crucial role in addressing distribution shifts in machine learning. In this work, we improve the theoretical foundations of UDA proposed in Acuna et al. (2021) by refining their $f$-divergence-based discrepancy and additionally introducing a new measure, $f$-domain discrepancy ($f$-DD). By removing the absolute value function and incorporating a scaling parameter, $f$-DD obtains novel target error and sample complexity bounds, allowing us to recover previous KL-based results and bridging the gap between algorithms and theory presented in Acuna et al. (2021). Using a localization technique, we also develop a fast-rate generalization bound. Empirical results demonstrate the superior performance of $f$-DD-based learning algorithms over previous works in popular UDA benchmarks.
@@ 1 Introduction
Machine learning often faces the daunting challenge of domain shift, where the distribution of data in the testing environment differs from that used in training. Unsupervised Domain Adaptation (UDA) arises as a solution to this problem. In UDA, models are allowed to access to labelled source domain data and unlabelled target domain data, while the ultimate goal is to find a model that performs well on the target domain. The mainstream theoretical foundations of UDA, and more broadly, domain adaptation [2], primarily rely on the seminal works of discrepancy-based theory [3, 4]. In particular, [3, 4] characterize the error gap between two domains using a hypothesis class-specified discrepancy measure e.g., $\mathcal{H}\Delta\mathcal{H}$-divergence. While these works initially focus on binary classification tasks and zero-one loss, [5] extend the theory to a more general setting. Subsequently, this kind of theoretical framework was extended by various works [6, 7, 8, 9, 10, 1, 11], all sharing some common properties such as the ability to estimate the proposed discrepancy from finite unlabeled samples. Importantly, these theoretical results often inspire the design of new algorithms, such as domain-adversarial training of neural networks (DANN) [12] and Margin Disparity Discrepancy (MDD) [9], directly motivated by $\mathcal{H}\Delta\mathcal{H}$-divergence.
@@
Recently, [1] proposes an $f$-divergence-based domain learning framework, which generalizes various previous frameworks (e.g., those based on $\mathcal{H}\Delta\mathcal{H}$-divergence) and have demonstrated great empirical successes. However, we argue that this framework has potential limitations, at least in three aspects.
@@
First, their discrepancy measure is based on the variational representation of $f$-divergence in [13] (cf. Lemma 2.1). Although this variational formula is commonly adopted, its weakness has been pointed out in several works [14, 15]. For example, from this formula, one cannot recover the well-known Donsker and Varadhan’s (DV) representation of KL divergence [16]. This reveals a second limitation: some existing domain adaptation or transfer learning theories are based on the DV representation of KL, such as [17, 18, 19, 20], and the framework by [1], although including KL as a special case of $f$-divergence, fail to unify those theories. Furthermore and more critically, the discrepancy measure proposed by [1] contains an absolute value function, while the original variational representation does not. Notably, this absolute value function is necessary in their derivation of a target error upper bound, and the fundamental reason behind this is still due to the weak version of the variational representation relied upon. Specifically, a variational representation of an $f$-divergence is a lower bound of the divergence, but using a weak lower bound creates technical difficulties in proving an upper bound for target domain error. [1] chooses to add an absolute value function accordingly, potentially leading to an overestimation of the corresponding $f$-divergence. In fact, this absolute value function is removed in their proposed algorithm, termed $f$-Domain Adversarial Learning ($f$-DAL). While $f$-DAL outperforms both DANN and MDD in standard benchmarks, this choice exhibits a clear gap between their theory and algorithm.
@@
In this work, to overcome these limitations and explore the full potential of $f$-divergence, we present an improved framework for $f$-divergence-based domain learning theory. Specifically, we apply a more advanced variational representation of $f$-divergence (cf. Lemma 2.2), independently developed by [21, 22] and [23]. After introducing some preliminaries, the organization of the remainder of the paper and our main contributions are summarized below.
@@
  • In Section 3, we revisit the theoretical analysis in [1], where we refine their $f$-divergence-based discrepancy by Lemma 2.2 while retaining the absolute value function in the definition. The resulting target error bound (cf. Lemma 3.1) and the KL-based generalization bound (cf. Theorem 3.1) complement the theoretical framework of [1].
@@
  • In Section 4, we design a novel $f$-divergence-based domain discrepancy measure, dubbed $f$-DD. Specifically, we eliminate the absolute value function from the definition used in Section 3, incorporating a scaling parameter instead. We then derive an upper bound for the target error (cf. Theorem 4.1) and the sample complexity bound for our $f$-DD. The generalization bound based on empirical $f$-DD naturally follows from these results. Notably, the obtained target error bound allows us to recover the previous KL-based result in [19] (cf. Corollary 4.1).
@@
  • In Section 5, to improve the convergence rate of our $f$-DD-based bound, we sharpen the bound using a localization technique [24, 10]. The localized $f$-DD allows for a crisp application of the local Rademacher complexity-based concentration results [24], while also enabling us to achieve a fast-rate target error bound (cf. Theorem 5.1). As a concrete example, we present a generalization bound based on localized KL-DD (cf. Theorem 5.2), where our proof techniques are directly connected to fast-rate PAC-Bayesian bounds [25, 26] and fast-rate information-theoretic generalization bounds [27, 28, 29].
@@
  • In Section 6, we conduct empirical studies based on our $f$-DD framework. We show $f$-DD outperforms the original $f$-DAL in three popular UDA benchmarks, with the best performance achieved by Jeffereys-DD. Additionally, we note that the training objective in $f$-DAL aligns more closely with our theory than with [1] (cf. Proposition 1). We also show that adding the absolute value function leads to overestimation and that optimizing the scaling parameter in $f$-DD may not be necessary in practical algorithms.
@@ 2 Preliminaries
Let $\mathcal{X}$ and $\mathcal{Y}$ be the input space and the label space. Let $\mathcal{H}$ be the hypothesis space, where each $h\in\mathcal{H}$ is a hypothesis mapping $\mathcal{X}$ to ${\cal Y}$. Consider a single-source domain adaptation setting, where $\mu$ and $\nu$ are two unknown distributions on $\mathcal{X}\times\mathcal{Y}$, characterizing respectively the source domain and the target domain. Let $\mathcal{S}=\{X_{i},Y_{i}\}_{i=1}^{n}{\sim}\mu^{\otimes n}$ be a labeled source-domain sample and $\mathcal{T}=\{X_{j}\}_{j=1}^{m}{\sim}\nu_{\cal X}^{\otimes m}$ be an unlabelled target-domain sample, where $\nu_{\cal X}$ is the marginal distribution of $X$ in the target domain. We use $\hat{\mu}$ and $\hat{\nu}$ to denote the empirical distributions on ${\cal X}$ corresponding to $\mathcal{S}$ and $\mathcal{T}$, respectively. The objective of UDA is to find a hypothesis $h\in\mathcal{H}$ based on $\mathcal{S}$ and $\mathcal{T}$ that “works well” on the target domain.
Let $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ be a symmetric loss (e.g., $\ell(y,y)=0$ for all $y\in{\cal Y}$). The population risk for each $h\in\mathcal{H}$ in the target domain (i.e. target error) is defined as $R_{\nu}(h)\triangleq{{\mathbb{E}_{(X,Y)\sim\nu}}\left[\ell(h(X),Y)\right]},$ and the population risk in the source domain, $R_{\mu}(h)$, is defined in the same way. Since $\mu$ and $\nu$ are unknown to the learner, one often uses recourse to the empirical risk in the source domain, which, for a given $\mathcal{S}$, is defined as $R_{\hat{\mu}}(h)\triangleq{{\mathbb{E}_{(X,Y)\sim\hat{\mu}}}\left[\ell(h(X),Y)\right]}={\frac{1}{n}\sum_{i=1}^{n}\ell(h(X_{i}),Y_{i})}$. Furthermore, let $f_{\nu}$ and $f_{\mu}$ be the ground truth labeling functions for the target domain and source domain, respectively, i.e. $f_{\nu}(x)=\arg\max_{y\in\mathcal{Y}}\nu(Y=y|x)$ and $f_{\mu}(x)=\arg\max_{y\in\mathcal{Y}}\mu(Y=y|x)$.
With a little abuse of the notation, we will simply use $\ell(h,h^{\prime})$ to represent $\ell(h(x),h^{\prime}(x))$ when the same $x$ is evaluated, serving as the “disagreement” for $h$ and $h^{\prime}$ on $x$. Additionally, following the conventional literature on DA theory [5], we assume that the loss function satisfies the triangle property [^1] The triangle property of loss function indicates that $\ell(y_{1},y_{2})\leq\ell(y_{1},y_{3})+\ell(y_{3},y_{2})$ for any $y_{1},y_{2},y_{3}\in\mathcal{Y}$.. For the readers’ convenience, a summary of all notations is provided in Table 5 in the Appendix.
@@
Definition 2.1 ($f$-divergence [30]).
Let $P$ and $Q$ be two distributions on $\Theta$. Let ϕ:+:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}$ be a convex function with $\phi(1)=0$. If $P\ll Q$ [^2] We say that $P$ is absolutely continuous with respect to $Q$, written $P\ll Q$, if $Q(A)=0\Longrightarrow P(A)=0$ for all measurable sets $A\subseteq\Theta$., then $f$-divergence is defined as ${\rm D}_{\phi}(P||Q)\triangleq{{\mathbb{E}_{Q}}\left[\phi\left(\frac{dP}{dQ}\right)\right]},$ where $\frac{dP}{dQ}$ is a Radon-Nikodym derivative.
The $f$-divergence family contains many popular divergences. For example, letting $\phi(x)=x\log{x}$ (or $x\log{x}+c(x-1)$ for any constant $c$) recovers the definition of KL divergence.
The $f$-divergence discrepancy measure by [1] is motivated by the variational formula of $f$-divergence that utilizes the Legendre transformation (LT).
Lemma 2.1 ([13]).
${\rm D}_{\phi}(P||Q)=\sup_{g\in\mathcal{G}}{{\mathbb{E}_{\theta\sim P}}\left[g(\theta)\right]}-{{\mathbb{E}_{\theta\sim Q}}\left[\phi^{*}(g(\theta))\right]}.$
However, it is well-known that the variational formula in Lemma 2.1 does not recover the the Donsker and Varadhan’s (DV) representation of KL divergence (cf. Lemma B.1). We will elaborate on this later. Recently, [23] and [21, 22] concurrently introduce a novel variational representation for $f$-divergence, which is also implicitly stated in [31, Theorem 4.2], as given below.
Lemma 2.2 ([21, Corollary 3.5]).
The variational formula of $f$-divergence is
This variational representation is a “shift transformation” of Lemma 2.1 (i.e. $g\to g+\alpha$). Note that this representation shares the same optimal solution as Lemma 2.1 (clearly identified as the corresponding $f$-divergence), but Lemma 2.2 is considered tighter in the sense that the representation in Lemma 2.2 is flatter around the optimal solution. [23] provides a comprehensive study to justify this, and they also show that Lemma 2.2 can accelerate numerical estimation of $f$-divergences.
Here, to illustrate the advantage of Lemma 2.2, we use KL divergence as an example. Specifically, let $\phi(x)=x\log{x}-x+1$, then its conjugate function is $\phi^{*}(y)=e^{y}-1$. Substituting $\phi^{*}$ into Lemma 2.1, we have
This representation is usually called LT-based KL. On the other hand, with the optimal $\alpha^{*}=-\log{{{{\mathbb{E}_{Q}}\left[e^{g(\theta)}\right]}}}$, Lemma 2.2 will give us
$\mathrm{D_{KL}}\left(P||Q\right)\!=$ $\sup_{g\in\mathcal{G}}{{\mathbb{E}_{P}}\left[g(\theta)\right]}\!\!-\!\inf_{\alpha\in\mathbb{R}}\left\{{{\mathbb{E}_{Q}}\left[e^{g(\theta)+\alpha}\right]}\!-\!1\!-\!\alpha\right\}\!=\sup_{g\in\mathcal{G}}{{\mathbb{E}_{P}}\left[g(\theta)\right]}\!-\!\log{{\mathbb{E}_{Q}}\left[e^{g(\theta)}\right]}.$ (2)
Notice that Eq. (2) immediately recovers the DV representation of KL. Since $\log(x)\leq x-1$ for $x>0$, it is evident that, as a lower bound of KL divergence, Eq. (2) is pointwise tighter than Eq. (1). In Appendix C, we also show the variational representations of some other divergences obtained from Lemma 2.2.
In the context of UDA, it may be tempting to think that using a point-wise smaller quantity (in Lemma 2.1), as the key component of an upper bound for target error, is essentially desired. However, neither Lemma 2.1 nor Lemma 2.2 is able to directly give such an upper bound. To elaborate, as $\phi^{*}(x)\geq x$ when $\phi(1)=0$ (cf. Lemma B.2), both Lemma 2.1 and Lemma 2.2 imply that $\mathrm{D}_{\phi}(P||Q)\leq\sup_{g}{{\mathbb{E}_{P}}\left[g(\theta)\right]}-{{\mathbb{E}_{Q}}\left[g(\theta)\right]}$. Bearing this in mind, UDA typically requires an upper bound for the quantity $\sup_{h}{{\mathbb{E}_{\nu}}\left[\ell\circ h(X)\right]}-{{\mathbb{E}_{\mu}}\left[\ell\circ h(X)\right]}$, and simply restricting $\mathcal{G}$ in Lemma 2.1 and Lemma 2.2 to a composition of $\mathcal{H}$ and $\ell$ can only provide a lower bound for such a quantity.
Before we propose an improved discrepancy measure, we first revisit the absolute discrepancy in [1] by using Lemma 2.2 instead. This also serves as a review of the common developments in the DA theory.
3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
Based on Lemma 2.2, we refine the discrepancy measure of [1, Definition 3] as follows.
@@ 3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
For a given $h\in\mathcal{H}$, the $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}$ discrepancy from $\mu$ to $\nu$ is defined as
@@
Removing the absolute value function in $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)$ does not alter its non-negativity. To see this, consider $h^{\prime}=h$. By the definitions of $\phi$ and $\phi^{*}$, $\inf_{\alpha}\phi^{*}(\alpha)-\alpha=\phi(1)=0$, we have ${{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}=I_{\phi,\nu}^{h}(\ell\circ h^{\prime})=0$. Consequently, since $h$ exists in $\mathcal{H}$, $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)\geq 0$ holds even without the absolute value function. In addition, due to the absolute value function, the relation between $\widetilde{\rm D}^{h,\mathcal{H}}_{\phi}(\mu||\nu)$ with the one in [1] is no longer clear.
@@
$R_{\nu}(h)\leq R_{\mu}(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+\lambda^{*}.$
Remark 3.2 (Regarding $\lambda^{*}$).
This error bound shares similarities with the previous works [3, 4, 9, 10, 1]. For example, the third term $\lambda^{*}$ is the ideal joint risk for the DA problem. As widely discussed in prior studies, this term captures the inherent challenge in the DA task and might be inevitable [4, 32]. However, it has also been pointed out that this $\lambda^{*}$ term can be significantly pessimistic, particularly in the case of conditional shift [33]. In fact, similar to the $\lambda^{*}$-free bound in [33, Theorem 4.1], it is a simple matter to replace $\lambda^{*}$ with the cross-domain error term $\min\{R_{\nu}(f_{\mu}),R_{\mu}(f_{\nu})\}$ in Lemma 3.1 (and all the other target error bounds in this paper). See Appendix D.3 for details. Given that [1] also uses $\lambda^{*}$, we use $\lambda^{*}$ in the bounds for consistency in the remainder of this paper.
In the sequel, we will give a Rademacher complexity based generalization bound for the target error. Let ^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ denote the empirical Rademacher complexity of function class =$\mathcal{F}=\{f:\mathcal{X}\to\mathbb{R}\}$ for some sample $S$ [34]. Notice that a shift transformation of a function class will not change its Rademacher complexity, so the generalization bound based on $\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}$ closely resembles the one presented in [1, Theorem 3], which contains a Lipschitz constant of $\phi^{*}$. Here, we give a generalization bound specialized for $\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}$, wherein the Lipschitz constant of $\phi^{*}$ can be explicitly determined.
@@
$R_{\nu}(h)\leq R_{\hat{\mu}}(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\hat{\mu}||\hat{\nu})+2e\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})+4\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})+{\lambda}^{*}+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}+\sqrt{\frac{\log(1/\delta)}{m}}\right),$